# When simulations look right but causal effects go wrong: Large language models as behavioral simulators

### 02 Apr 2026 v1

@@ Results
At baseline, all three models recovered broad cross-national descriptive patterns, but performance varied strongly by outcome domain (Fig. 2a-d). Country-level mean error was modest for belief (0.24-0.65 percentage points, p.p.) and policy (-0.55 to 8.63 p.p.), but much larger for action, where GPT and Gemini overpredicted action by about 13.6 and 7.9 p.p., while Claude underpredicted it by 12.8 p.p. (Fig. 2a). Individual-level mean absolute error (MAE) showed similar patterns, remaining  16-18 points for belief and policy but rising to  40-43 points for action (Fig. 2b). Country-level structural alignment was also modest, with baseline Spearman correlations of 0.29-0.39 and KNN overlap of the most similar five countries of 0.21-0.25 (Fig. 2c). We also conducted baseline comparisons using multiple linear regression and LASSO (Supplementary Table S1), suggesting that LLMs could approach the individual-level descriptive performance of conventional statistical baselines despite operating without training examples.
@@
Prompt enhancements further strengthened this descriptive fit. From a broader set of candidates spanning external information provision and internal reasoning elicitation (Fig. 2e; Supplementary Table S2), we selected two best-performing methods using the control group as a development sample (n = 5,093): KNN-based few-shot prompting (few-shot) and VBN-based chain-of-thought reasoning (VBN-CoT). Both improved descriptive fit. Country-level mean error decreased by 0.6-5.6 p.p. depending on outcome and method, and individual-level MAE gains were 0.5-1.4 p.p. (Fig. 2f-g). Few-shot produced larger gains in country-level alignment than VBN-CoT and raised Spearman correlation from about 0.39 to 0.65 (Fig. 2h). VBN-CoT produced larger reductions in distributional divergence for belief and policy, although neither method substantially improved distributional fit for action (Fig. 2i). At the same time, all models and prompting enhancements used exhibited variance compression, with predicted variance consistently below observed variance across all models and outcomes (Fig. 2d), indicating that descriptive fit was strongest for central tendencies rather than the full spread of human responses.
@@
However, this descriptive fit did not carry over to causal fidelity. Under baseline prompting, when averaged across outcomes, mean absolute ATE error was 9.5 p.p. for GPT, 8.8 p.p. for Gemini, and 9.2 p.p. for Claude (Fig. 2j). These errors were especially larger for action, consistent with the descriptive results above. By contrast, linear regression and LASSO had lower mean absolute ATE errors of 4.3 p.p. and 4.2 p.p., respectively (Supplementary Table S1), suggesting that the limited causal fidelity of the LLM simulations was not solely a consequence of task difficulty.
@@
Prompt refinements improved this picture selectively. VBN-CoT reduced mean absolute ATE error by 34.0%-41.2% relative to baseline across all three models, but these gains were concentrated in belief and policy (e.g., 12.5 p.p. and 6.5 p.p. to 3.9 p.p. and 3.5 p.p. for GPT), with little benefit for action (from 8.4 p.p. to 8.7 p.p.). Few-shot was highly model-dependent and in some cases worsened causal error for action (e.g., from 8.4 p.p. to 44.2 p.p. in Claude and from 16.4 p.p. to 36.7 p.p. in Gemini). In neither dataset did the prompting strategy that minimized individual-level MAE also minimize ATE error (Supplementary Table S3). Thus, the same prompting strategies that made the simulations look more descriptively reliable did not produce parallel improvements in causal fidelity.
@@
The divergence between descriptive fit and causal fidelity also represents a difference in error structure (Fig. 3a). In descriptive evaluation, error variance was dominated by outcome type among all three models (93.3%-94.5%), and prompting method contributed less than 1%. However, the structure of causal error reorganized. The contribution of outcome type declined to 11.9%-40.8%, while method and outcome-by-method interactions jointly accounted for 24.1%-65.1% of variance. The three models showed similar qualitative shifts from outcome-dominated descriptive error to a more distributed causal error structure, although the dimensions most amplified differed somewhat across models. GPT showed greater country-related heterogeneity (49.9% from country and its interactions), Claude showed the strongest amplification of method and its interactions (65.1% combined), and Gemini retained both a relatively large outcome main effect (40.8%) and substantial method-related variance (40.2% combined for method and outcome-by-method interactions).
Similar structural shifts replicated in two additional datasets (Fig. 3b). Variance decomposition likewise showed that descriptive and causal error diverged along different dimensions in the two designs. Method played a larger role in causal than descriptive error (15.8% vs. 3.2%) in the Spampatti dataset, while in the Većkalov dataset the divergence was carried more by shifts in the relative contributions of outcome (from 2.7% to 28.4%) and country (from 64.1% to 49.7%). Across datasets, the divergence was robust, although the dimensions along which it emerged varied with experimental design.
Precision-weighted and multilevel analyses further specified this structural difference. Intraclass correlation coefficients (ICC) for country random intercepts increased from 0.8%-1.6% for MAE to 13.1%-16.1% for absolute ATE error (Fig. 3c), indicating substantially greater between-country heterogeneity in causal evaluations. ICCs were amplified in causal evaluations in both replication datasets (Spampatti from 0.3% to 29.2%, Većkalov from 4.9% to 53.9%). Permutation tests shuffling responses within country-by-outcome blocks further showed the structural difference between MAE and ATE error was unlikely to arise from block composition alone, in all three datasets (all p < 0.001). Therefore, descriptive fit and causal fidelity were shaped by different sources of error, so descriptive evaluation alone would miss some of the most consequential sources of causal error.
@@
The contrast between attitudinal and behavioral outcomes became clearer when we changed intervention content and outcome framing separately (Fig. 4c-d). First, three representative interventions were edited to vary persuasive force while preserving overall structure: (1) surface paraphrase, (2) force removal, and (3) force reversal (Supplementary Tables LABEL:tab:s5a,  LABEL:tab:s5b, and  LABEL:tab:s5c). Force removal reduced the ATE overestimation for belief and policy, with the largest reductions for scientific consensus (belief: -4.49, policy: -3.68), intermediate reductions for dynamic norm (-2.32, -1.76), and the smallest reductions for future-self continuity (-1.15, -0.26). But ATE error for action changed less coherently under the same perturbations. Then, we altered how the attitude-behavior relation was framed without changing intervention content (Supplementary Table LABEL:tab:s5d). Instructing the model to infer attitudes and behaviors separately produced little change in belief or policy estimates and did not improve action ATE error consistently across interventions. Explicitly stating that behavioral change is typically smaller and slower than attitudinal change shifted action estimates, reducing ATE error for scientific consensus by 0.16 (from 1.36 to 1.20), leaving it largely unchanged for dynamic norm (-0.01, from 0.10 to 0.09), and overcorrecting for future-self continuity (+0.64, from 0.24 to 0.88, with bias flipping from overestimation to underestimation). These probes suggest that attitudinal simulations responded more consistently than behavioral ones to intervention content, and behavioral simulations were more sensitive to how the attitude-behavior coupling was framed. Therefore, causal distortion was especially marked when the simulation required linking attitudinal and behavioral domains.
@@
Figure 5: Countries and population groups well captured descriptively are not necessarily those with accurate causal estimates. Results shown for GPT; other models in Supplementary Information. a. Descriptive and causal error are weakly correlated at the country level. Open markers indicate baseline, and filled markers indicate prompting refinements. b. Countries with highest and lowest descriptive error. c. Countries with highest and lowest causal error. Overlap between b and c is minimal. d-e. Country-level descriptive gains concentrate in OECD and high-internet countries, but these advantages do not carry over to causal fidelity. Error bars show 95% CI. f-g. Descriptive disparities are largest along age and gender, while causal disparities are largest along political orientation and SES. Error bars show 95% CI.
@@
At the country level, descriptive and causal error were only weakly associated (Fig. 5a). Baseline MAE and ATE error were only weakly associated among all three outcomes (belief r = 0.22; policy r = 0.12; action r = -0.08). This divergence was also visible in country rankings (Fig. 5b–c). The 10 lowest-error countries for MAE and ATE error shared only 2 members (Saudi Arabia and Venezuela), and the 10 highest-error countries overlapped on only 3 (Italy, China, and Portugal). In other words, a country that looked descriptively well captured was not necessarily one in which intervention effects were causally well recovered. Relying solely on descriptive evaluation would fail to identify the countries with the largest causal errors.
The same divergence changed which populations appeared most advantaged or most at risk under prompting improvements (Fig.  5d-g). At the country level, descriptive gains were modestly concentrated in OECD countries and countries with higher internet penetration, with between-group differences typically within 1-3 p.p. (e.g., VBN-CoT, OECD vs. non-OECD: 1.41 p.p. larger, p = 0.005; high vs. low internet: 1.99 p.p. greater, p = 0.002). On the causal side, these structures changed as between-group differences in ATE error were unstable and shifted direction across methods. OECD countries showed slightly higher causal gains under few-shot (difference: 0.68, p = 0.069) but lower gains under VBN-CoT (-0.75, p = 0.199). Country characteristics that appeared mildly advantageous at the descriptive level did not reliably translate into causal-level advantages. This divergence was more pronounced at the demographic level. Descriptive gains were largest along age and gender dimensions (e.g., 18-24 vs. 65+ age groups: 4.48 p.p. under few-shot and 5.28 p.p. under VBN-CoT; female groups consistently showed  1 p.p. greater MAE improvement than the male group; both p < 0.001). By contrast, the most prominent causal disparities appeared along political orientation (e.g. right-leaning vs. left-leaning: 7.74 p.p. under few-shot and 5.76 under VBN-CoT; p < 0.001) and SES (high vs. low: 3.55 and 2.92 p.p. for the two methods, p < 0.001), while gender and age differences were not significant. The inequality dimensions most salient in descriptive evaluation did not retain their prominence in causal fidelity.
@@
Figure 6: Exploratory feature importance profiles for descriptive and causal error. Descriptive error changes are primarily associated with statistical prediction difficulty (e.g., true mean, original MAE), while causal error changes are more strongly linked to socioeconomic features (e.g., Gini coefficient, internet penetration). Signed SHAP percentage values from random forest models (100-seed averaged) predicting enhancement-induced changes in MAE and ATE error across outcomes. a. Few-shot. b. VBN-CoT. Positive values (red) indicate that higher feature values are associated with larger error increases; negative values (blue) indicate larger error reductions. Values were defined relative to the baseline prompt.
An exploratory SHAP analysis further indicated that enhancement-induced changes in descriptive and causal error were associated with different feature profiles (Fig. 6). Changes in MAE were primarily associated with statistical features. The normalized signed SHAP contribution of true mean to predicted changes in action MAE reached 74.6%, with original MAE and median age also contributing substantially. By contrast, changes in ATE error were more strongly associated with socioeconomic features. For instance, the SHAP value of Gini coefficient for belief ATE error was 31.6% (few-shot) and 25.0% (VBN-CoT), and the contributions of internet penetration rate and political orientation were also markedly elevated, although directions varied across outcomes and enhancement methods. Therefore, enhancement-induced changes in descriptive error were more closely linked to statistical prediction difficulty, whereas changes in causal error were more strongly associated with mismatches between national social structures and the model’s assumptions about how attitudes translate into behavior.
@@ Discussion
Populations and societies that are well represented descriptively are not necessarily those whose responses to intervention are inferred more accurately. Our results show that the inequality structures at the descriptive and causal levels follow different patterns. Descriptive improvements were more concentrated along demographic patterns such as age and gender, while causal bias fell more heavily on political orientation and socioeconomic status. A similar mismatch appeared at the country level. Descriptive improvements were more concentrated in OECD countries and in countries with higher internet penetration, but these advantages did not carry over consistently to causal fidelity. Moreover, prompting strategies introduced to improve simulation quality also shift the distribution of error across groups and settings. Therefore, a group or society may be closely matched to real data in means and distributions and still have its response to intervention misread. Fairness here concerns both who is represented more accurately and whose mechanisms of change are better understood. If descriptive validation remains the primary standard, current fairness frameworks may not be sufficient to surface the real distributional risks of these systems.
@@
Country-level inequality. To assess whether cross-national reliability gaps change when moving from descriptive to causal fidelity, we stratified countries by four structural characteristics: OECD membership (binary, to capture economic development), internet penetration (tercile split, to capture digital access), English proficiency (tercile split, to capture potential linguistic alignment with model training data), and income inequality (Gini coefficient, tercile split, to capture socio-economic conditions). For each stratification, between-group comparisons focused on the extreme groups (e.g., High vs. Low for English proficiency, top vs. bottom tercile for internet penetration and Gini). We compared between-group differences in mean MAE / mean ATE error and estimated 95% bootstrap confidence intervals based on 5,000 resamples. We additionally used Welch’s t-test and Mann-Whitney U test to assess the robustness of group differences. Except for English proficiency scores, which were obtained from the EF English Proficiency Index (EF EPI), all other country-level data were collected from Our World in Data.
@@ Acknowledgements
   Action 39.75 41.50 40.63 42.13 43.13
Absolute ATE error 4.32 4.24 9.47 8.85 9.23
   Belief 2.99 2.97 7.83 6.52 12.53
   Policy 2.87 2.85 5.93 3.42 6.52
   Action 7.10 6.90 14.65 16.59 8.64
@@
Spampatti VBN-CoT 29.70 3 4.30 1
Većkalov Baseline 10.61 2 2.04 2
Većkalov Few-shot 10.60 1 2.07 3
Većkalov VBN-CoT 12.01 3 2.00 1
@@
Table S10: Results of Welch’s $t$-test for group comparisons (MAE). ∗∗∗ $p<0.001$, ∗∗ $p<0.01$, $p<0.05$ when both tests were significant; indicates that only one of the two tests reached $p<0.05$; unmarked cells were not significant in either test.
@@
Gini High vs. Low 0.76-0.76 +0.48+0.48 $4.91^{**}$ 0.58-0.58 $2.96^{*}$ +0.76+0.76
Internet rate High vs. Low +0.57+0.57 0.92-0.92 $3.37^{*} +$$1.51^{*} +$1.72+1.72 $1.28^{*}$
English proficiency High vs. Low 0.15-0.15 0.39-0.39 $5.13^{**} +$$1.97^{**} +$1.55+1.55 $1.02^{\dagger}$
OECD OECD vs. Non-OECD +0.68+0.68 0.75-0.75 $3.39^{**} +$$26^{*} +1.1$9+1.19 $1.20^{**}$
Age 18–24 vs. 65+ +1.06+1.06 +0.97+0.97 $4.84^{***}$ $0.91^{\dagger}$ $2.44^{*} +$0.14+0.14
Gender Male vs. Female 0.12-0.12 0.36-0.36 +0.45+0.45 +0.15+0.15 $1.22^{*} +$0.24+0.24
Political orientation Left vs. Right $7.74^{***}$ $5.76^{***}$ $2.84^{**} +$0.27+0.27 $3.89^{***}$ $5.48^{***}$
SES High vs. Low $3.55^{***} +$$2.92^{***}$ $0.88^{\dagger}$ $1.05^{*}$ $1.66^{*}$ 0.02-0.02
Table S11: Results of Welch’s $t$-test for group comparisons (ATE). ∗∗∗ $p<0.001$, ∗∗ $p<0.01$, $p<0.05$ when both tests were significant; indicates that only one of the two tests reached $p<0.05$; unmarked cells were not significant in either test.
@@
Gini High vs. Low 0.05-0.05 0.15-0.15 0.21-0.21 $2.55^{**}$ 0.10-0.10 $2.32^{***}$
Internet rate High vs. Low $1.25^{*} +$$1.99^{**} +$0.06+0.06 +1.24+1.24 +1.01+1.01 $2.66^{***}$
English proficiency High vs. Low +0.27+0.27 +1.09+1.09 +0.13+0.13 $2.14^{*} +$0.27+0.27 $2.42^{***}$
OECD OECD vs. Non-OECD +0.58+0.58 $1.41^{**}$ 0.16-0.16 $1.32^{*} +$0.62+0.62 $2.20^{***}$
Age 18–24 vs. 65+ $4.48^{***}$ $5.28^{***}$ $2.56^{***}$ $6.67^{***}$ $4.98^{***}$ $0.57^{\dagger}$
Gender Male vs. Female $0.98^{***}$ $1.10^{***}$ $0.66^{*}$ $1.63^{***}$ $1.01^{***}$ $0.62^{**}$
Political orientation Left vs. Right $1.90^{***}$ +0.11+0.11 $0.97^{*} +$$2.09^{***} +$0.36+0.36 $1.58^{***}$
SES High vs. Low +0.06+0.06 +0.41+0.41 $2^{\dagger} +0.77+0$.77 $0.78^{*}$ $0.92^{**}$

### 10 Apr 2026 v2

@@ Results
At baseline, all three models recovered broad cross-national descriptive patterns, but performance varied strongly by outcome domain (Fig. 2a-d). Country-level mean error was modest for belief (0.24-6.65 percentage points, p.p.) and policy (-0.55 to 8.63 p.p.), but much larger for action, where GPT and Gemini overpredicted action by about 13.6 and 7.9 p.p., while Claude underpredicted it by 12.8 p.p. (Fig. 2a). Individual-level mean absolute error (MAE) showed similar patterns, remaining  16-18 points for belief and policy but rising to  40-43 points for action (Fig. 2b). Country-level structural alignment was also modest, with baseline Spearman correlations of 0.29-0.39 and KNN overlap of the most similar five countries of 0.17-0.23 (Fig. 2c). We also conducted baseline comparisons using multiple linear regression and LASSO (Supplementary Table S1), suggesting that LLMs could approach the individual-level descriptive performance of conventional statistical baselines despite operating without training examples.
@@
Prompt enhancements further strengthened this descriptive fit. From a broader set of candidates spanning external information provision and internal reasoning elicitation (Fig. 2e; Supplementary Table S2), we selected two best-performing methods using the control group as a development sample (n = 5,093): KNN-based few-shot prompting (few-shot) and VBN-based chain-of-thought reasoning (VBN-CoT). Both generally improved descriptive fit. For GPT, country-level mean error decreased by 0.6-5.6 p.p. depending on outcome and method, and individual-level MAE gains were typically around 0.2-1.4 p.p. (Fig. 2f-g), although not uniformly across outcomes. Few-shot produced larger gains in country-level alignment than VBN-CoT and raised Spearman correlation from about 0.39 to 0.65 (Fig. 2h). VBN-CoT produced larger reductions in distributional divergence for belief and policy, although neither method substantially improved distributional fit for action (Fig. 2i). At the same time, all models and prompting enhancements used exhibited variance compression, with predicted variance consistently below observed variance across all models and outcomes (Fig. 2d), indicating that descriptive fit was strongest for central tendencies rather than the full spread of human responses.
@@
However, this descriptive fit did not carry over to causal fidelity. Under baseline prompting, when averaged across outcomes, mean absolute ATE error was 9.4 p.p. for GPT, 8.7 p.p. for Gemini, and 9.1 p.p. for Claude (Fig. 2j). These errors were especially larger for action, consistent with the descriptive results above. By contrast, linear regression and LASSO had lower mean absolute ATE errors of 4.3 p.p. and 4.2 p.p., respectively (Supplementary Table S1), suggesting that the limited causal fidelity of the LLM simulations was not solely a consequence of task difficulty.
@@
Prompt refinements improved this picture selectively. VBN-CoT reduced mean absolute ATE error by 33.9%-42.8% relative to baseline across all three models, with 3.9 p.p. for belief, 3.5 p.p. for policy, and 8.7 p.p. for action, compared with 7.8, 5.9, and 14.5 p.p. under baseline prompting for GPT. Few-shot was highly model-dependent and in some cases worsened causal error for action (e.g., from 8.4 p.p. to 44.2 p.p. in Claude and from 16.4 p.p. to 36.7 p.p. in Gemini). In neither dataset did the prompting strategy that minimized individual-level MAE also minimize ATE error (Supplementary Table S3). Thus, the same prompting strategies that made the simulations look more descriptively reliable did not produce parallel improvements in causal fidelity.
@@
The divergence between descriptive fit and causal fidelity also represents a difference in error structure (Fig. 3a). In descriptive evaluation, error variance was dominated by outcome type among all three models (92.8%-94.5%), and prompting method contributed less than 1%. However, the structure of causal error reorganized. The contribution of outcome type declined to 20.6%-40.8%, while method and outcome-by-method interactions jointly accounted for 21.4%-65.1% of variance. The three models showed similar qualitative shifts from outcome-dominated descriptive error to a more distributed causal error structure, although the dimensions most amplified differed somewhat across models. GPT showed greater country-related heterogeneity (41.8% from country and its interactions), Claude showed the strongest amplification of method and its interactions (65.1% combined), and Gemini retained both a relatively large outcome main effect (40.8%) and substantial method-related variance (40.2% combined for method and outcome-by-method interactions).
Similar structural shifts replicated in two additional datasets (Fig. 3b). Variance decomposition likewise showed that descriptive and causal error diverged along different dimensions in the two designs. In the Spampatti dataset, this divergence reflected a reallocation of variance from the outcome-method interaction to the main effects of outcome and method, with the latter contributing more strongly to causal error. While in the Većkalov dataset, the divergence was carried more by shifts in the relative contributions of outcome (from 2.7% to 28.4%) and country (from 64.1% to 49.7%). Across datasets, the divergence was robust, although the dimensions along which it emerged varied with experimental design.
Precision-weighted and multilevel analyses further specified this structural difference. Intraclass correlation coefficients (ICC) for country random intercepts increased from 0.8%-1.6% for MAE to 11.8%-16.1% for absolute ATE error (Fig. 3c), indicating substantially greater between-country heterogeneity in causal evaluations. ICCs were amplified in causal evaluations in both replication datasets (Spampatti from 0.3% to 29.2%, Većkalov from 4.9% to 53.9%). Permutation tests shuffling responses within country-by-outcome blocks further showed the structural difference between MAE and ATE error was unlikely to arise from block composition alone, in all three datasets (all p < 0.001). Therefore, descriptive fit and causal fidelity were shaped by different sources of error, so descriptive evaluation alone would miss some of the most consequential sources of causal error.
@@
The contrast between attitudinal and behavioral outcomes became clearer when we changed intervention content and outcome framing separately (Fig. 4c-d). First, three representative interventions were edited to vary persuasive force while preserving overall structure: (1) surface paraphrase, (2) force removal, and (3) force reversal (Supplementary Tables LABEL:tab:s5a,  LABEL:tab:s5b, and  LABEL:tab:s5c). Force reversal reduced the ATE overestimation for belief and policy, with the largest reductions for scientific consensus (belief: -4.49, policy: -3.71), intermediate reductions for dynamic norm (-2.37, -2.15), and the smallest reductions for future-self continuity (-1.13, -0.22). But ATE error for action changed less coherently under the same perturbations. Then, we altered how the attitude-behavior relation was framed without changing intervention content (Supplementary Table LABEL:tab:s5d). Instructing the model to infer attitudes and behaviors separately produced little change in belief or policy estimates and did not improve action ATE error consistently across interventions. Explicitly stating that behavioral change is typically smaller and slower than attitudinal change shifted action estimates, reducing ATE error for scientific consensus by 0.16 (from 1.36 to 1.20), leaving it largely unchanged for dynamic norm (-0.01, from 0.10 to 0.09), and overcorrecting for future-self continuity (+0.64, from 0.24 to 0.88, with bias flipping from overestimation to underestimation). These probes suggest that attitudinal simulations responded more consistently than behavioral ones to intervention content, and behavioral simulations were more sensitive to how the attitude-behavior coupling was framed. Therefore, causal distortion was especially marked when the simulation required linking attitudinal and behavioral domains.
@@
Figure 5: Countries and population groups well captured descriptively are not necessarily those with accurate causal estimates. Results shown for GPT; other models in Supplementary Information. a. Descriptive and causal error are weakly correlated at the country level. Open markers indicate baseline, and filled markers indicate prompting refinements. b. Countries with highest and lowest descriptive error. c. Countries with highest and lowest causal error. Overlap between b and c is minimal. d-e. Country-level descriptive gains concentrate in OECD, high-internet, English-proficient, and low-inequality countries, but these advantages do not carry over to causal fidelity. Error bars show 95% CI. f-g. Demographic moderators are weaker and less consistent than country-level characteristics for both descriptive and causal gains. Error bars show 95% CI.
@@
At the country level, descriptive and causal error were only weakly associated (Fig. 5a). Baseline MAE and ATE error were only weakly associated among all three outcomes (belief r = 0.22; policy r = 0.15; action r = 0.11). This divergence was also visible in country rankings (Fig. 5b–c). The 10 lowest-error countries for MAE and ATE error shared only 3 members (Chile, China, and Portugal), and the 10 highest-error countries overlapped on only 1 (Morocco). In other words, a country that looked descriptively well captured was not necessarily one in which intervention effects were causally well recovered. Relying solely on descriptive evaluation would fail to identify the countries with the largest causal errors.
The distribution of prompting refinement gains across populations also differed between descriptive and causal levels (Fig.  5d-g). At the country level, descriptive gains were systematically concentrated in more connected and economically equal countries with larger improvements in OECD countries and in countries with higher internet penetration, higher English proficiency, and lower Gini inequality (0.6 to 2.0 p.p.; all adjusted p < .05 for both methods). Under VBN-CoT, political orientation showed the largest descriptive gap (+2.08 p.p., adjusted p < .001), and age also showed a smaller but significant difference (18-24 vs. 65+: -0.24 p.p., adjusted p = .014). On the causal side, this structured inequality largely dissolved. No analyzed country-level characteristics were significantly associated with ATE improvement under either refinement. Political orientation and SES showed notable associations under VBN-CoT (political orientation: +1.62, p < .001; SES: -1.42, p < .01) but not under few-shot. Therefore, prompting refinements produced a clear and method-consistent descriptive pattern across countries, but the inequality dimensions most salient in descriptive evaluation did not retain their prominence in causal fidelity.
@@
Figure 6: Exploratory feature importance profiles for descriptive and causal error. Descriptive error changes are largely associated with statistical prediction difficulty (e.g., true mean, original MAE), whereas causal error changes show a more heterogeneous mix of statistical, demographic, and country-level features. Signed SHAP percentage values from random forest models (100-seed averaged) predicting enhancement-induced changes in MAE and ATE error across outcomes. a. Few-shot. b. VBN-CoT. Positive values (red) indicate that higher feature values are associated with larger error increases; negative values (blue) indicate larger error reductions. Values were defined relative to the baseline prompt.
An exploratory SHAP analysis further indicated that changes induced by prompting refinements in descriptive and causal error were associated with different feature profiles (Fig. 6). Changes in MAE were largely associated with statistical features. The country’s true response mean and baseline prediction difficulty together dominated the feature profile, with individual contributions reaching 42.1% (true mean for VBN-CoT belief) and 41.0% (original MAE for few-shot policy). By contrast, changes in ATE error showed a more heterogeneous pattern across statistical, demographic, and country-level features, and the direction of associations frequently shifted across outcome domains and prompting methods. Descriptive gains thus tracked predictable statistical properties of the simulation task, whereas causal gains were shaped by task-specific interactions that no single country characteristic reliably captured.
@@ Discussion
Populations and societies that are well represented descriptively are not necessarily those whose responses to intervention are inferred more accurately. Our results show that the inequality structures at the descriptive and causal levels follow different patterns. Descriptive improvements were more concentrated in OECD countries and in countries with higher internet penetration and English proficiency, whereas comparable structure was largely absent for causal fidelity. Moreover, prompting strategies introduced to improve simulation quality also shift the distribution of error across groups and settings. Therefore, a group or society may be closely matched to real data in means and distributions and still have its response to intervention misread. Fairness here concerns both who is represented more accurately and whose mechanisms of change are better understood. If descriptive validation remains the primary standard, current fairness frameworks may not be sufficient to surface the real distributional risks of these systems.
@@
Country-level inequality. To assess whether cross-national reliability gaps change when moving from descriptive to causal fidelity, we stratified countries by four structural characteristics: OECD membership (binary, to capture economic development), internet penetration (tercile split, to capture digital access), English proficiency (tercile split, to capture potential linguistic alignment with model training data), and income inequality (Gini coefficient, tercile split, to capture socio-economic conditions). For each stratification, between-group comparisons focused on the extreme groups (e.g., High vs. Low for English proficiency, top vs. bottom tercile for internet penetration and Gini). We compared between-group differences in mean MAE / mean ATE error and estimated 95% bootstrap confidence intervals based on 5,000 resamples. We used pairwise Mann-Whitney U tests with Benjamini-Hochberg correction for multiple comparisons to assess subgroup differences. Except for English proficiency scores, which were obtained from the EF English Proficiency Index (EF EPI), all other country-level data were collected from Our World in Data.
@@
Data availability
The datasets analyzed in this study are publicly available from the original publications cited in the main text. Information on data sources and access is provided in the corresponding articles.
Code availability
Code used for analysis is available on Github.
@@ Acknowledgements
   Action 39.75 41.50 40.65 42.13 43.15
Absolute ATE error 4.32 4.24 9.39 8.72 9.14
   Belief 2.99 2.97 7.80 6.47 12.51
   Policy 2.87 2.85 5.86 3.34 6.47
   Action 7.10 6.90 14.51 16.35 8.43
@@
Spampatti VBN-CoT 29.70 3 4.31 1
Većkalov Baseline 10.74 2 3.20 2
Većkalov Few-shot 10.73 1 3.22 3
Većkalov VBN-CoT 11.51 3 2.91 1
@@
Table S10: Pairwise group comparisons for MAE. Entries report group differences based on pairwise Mann-Whitney U tests, with significance markers defined by BH-adjusted $p$-values ($p_{\mathrm{adj}}$): ∗∗∗ $p_{\mathrm{adj}}<0.001$, ∗∗ $p_{\mathrm{adj}}<0.01$, $p_{\mathrm{adj}}<0.05$, $p_{\mathrm{adj}}<0.1$. Unmarked cells were not significant after BH correction.
@@
Gini High vs. Low $^{***} 1.64-1.64^{**}$ 0.18-0.18 $2.82^{**}$ +0.05+0.05 $2.38^{***}$
Internet rate High vs. Low $1.00^{**} +$$1.95^{***} +$0.10+0.10 $1.53^{*} +$0.87+0.87 $2.74^{***}$
English proficiency High vs. Low $0.67^{*} +$$1.68^{***} +$0.22+0.22 $2.10^{*} +$0.32+0.32 $2.48^{***}$
OECD OECD vs. Non-OECD $0.64^{**} +$$1.62^{***}$ 0.12-0.12 $1.51^{*} +$0.54+0.54 $2.27^{***}$
Age 18–24 vs. 65+ +0.49+0.49 $0.24^{*}$ $2.62^{***}$ $6.66^{***}$ $5.03^{***}$ $0.58^{*}$
Gender Male vs. Female +0.11+0.11 0.07-0.07 $0.60^{*}$ $1.67^{***}$ $0.98^{***}$ $0.63^{***}$
Political orientation Left vs. Right +0.12+0.12 $2.08^{***} +$$0.93^{***} +$$2.09^{***} +$0.36+0.36 $1.57^{***}$
SES High vs. Low +0.13+0.13 $0.42^{\dagger}$ $0.67^{***}$ +0.80+0.80 $0.73^{***}$ $0.92^{***}$
Table S11: Pairwise group comparisons for ATE. Entries report group differences based on pairwise Mann-Whitney U tests, with significance markers defined by BH-adjusted $p$-values ($p_{\mathrm{adj}}$): ∗∗∗ $p_{\mathrm{adj}}<0.001$, ∗∗ $p_{\mathrm{adj}}<0.01$, $p_{\mathrm{adj}}<0.05$, $p_{\mathrm{adj}}<0.1$. Unmarked cells were not significant after BH correction.
@@
Gini High vs. Low $0.76^{\dagger}$ +0.48+0.48 $4.91^{**}$ 0.58-0.58 $2.96^{*}$ $0.76^{\dagger}$
Internet rate High vs. Low +0.33+0.33 0.94-0.94 $3.14^{*} +$$1.66^{*} +$1.33+1.33 $1.36^{*}$
English proficiency High vs. Low 0.44-0.44 0.46-0.46 $4.77^{*} +$$2.15^{**} +$1.07+1.07 $1.13^{*}$
OECD OECD vs. Non-OECD +0.51+0.51 0.77-0.77 $3.22^{**} +$$1.37^{**} +$0.92+0.92 $1.27^{**}$
Age 18–24 vs. 65+ 0.46-0.46 0.59-0.59 $4.85^{***}$ $0.93^{*}$ $2.47^{***} +$0.07+0.07
Gender Male vs. Female +0.28+0.28 +0.02+0.02 +0.49+0.49 +0.14+0.14 $1.27^{**} +$0.22+0.22
Political orientation Left vs. Right 0.39-0.39 $1.62^{***} +$$2.87^{**} +$0.26+0.26 $3.97^{***}$ $5.51^{***}$
SES High vs. Low 0.72-0.72 $1.42^{**}$ $0.85^{\dagger}$ $1.04^{*}$ $1.64^{***}$ 0.01-0.01